The fundamental frequency of an open organ pipe is 261.6Hz. The third resonance of a closed organ pipe has the same frequency. If the speed of sound in air is 340m/s, what is the length of the pipes? 




1
Expert's answer
2021-02-22T10:27:36-0500

Let's first find the wavelength of the wave from the wave speed equation:


"v=f\lambda,""\lambda=\dfrac{v}{f}=\dfrac{340\ \dfrac{m}{s}}{261.6\ Hz}=1.3\ m."

a) For the fundamental frequency (or the first harmonic) of an open organ pipe there is the following length-wavelength relationship:


"\lambda=2L,""L=\dfrac{\lambda}{2}=\dfrac{1.3\ m}{2}=0.65\ m=65\ cm."

b) For the third resonance frequency (or the third harmonic) of the closed organ pipe there is the following length-wavelength relationship:


"\lambda=\dfrac{4}{3}L,""L=\dfrac{3}{4}\lambda=\dfrac{3}{4}\cdot1.3\ m=0.975\ m=97.5\ cm."